OpenStudy (anonymous):

exact values for secx=-2/sqrt3

OpenStudy (mimi_x3):

\[secx = -\frac{2}{\sqrt{3}} \] \[secx = \frac{1}{cosx} \]

OpenStudy (mimi_x3):

I think that you're able to do it now..

OpenStudy (mimi_x3):

\[\frac{1}{cosx} = -\frac{2}{\sqrt{3}} \]

OpenStudy (anonymous):

no you'll have to explain a bit better than that, i can tell you the answer is 5pi/6 and -5pi/6 but im not sure how to get there

OpenStudy (mimi_x3):

well cross multiply

OpenStudy (mimi_x3):

\[\frac{1}{cosx} = \frac{-2}{\sqrt{3}} => \sqrt{3} = -2cosx\] you should be able to do it now :)

OpenStudy (anonymous):

thanks ill keep trying, would i b using the fourth quadrant to draw this

OpenStudy (mimi_x3):

 since cosx is a negative; so you will use the quadrants that i marks

OpenStudy (anonymous):

that getting clearer for me, now if you could just name the angles with the answers i gave before i might be able to complete the puzzle

OpenStudy (anonymous):

i think the angle has to be in one quadrant or the other

OpenStudy (anonymous):

oops i gave you the wrong answer a moment ago, the correct answer is pi/3 and -2pi/3, maybe that will help

OpenStudy (mimi_x3):

hmm; i dont know how to explain sorry \[ cosx = -\frac{\sqrt{3}}{2} => x = \frac{\pi}{6} + \pi => \frac{7\pi}{6} \]

OpenStudy (mimi_x3):

 \[ x = \pi - \frac{\pi}{6} = \frac{5\pi}{6}\]